OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of a double angle. Then find the exact value of the expression. 2 sin 120° cos 120°

OpenStudy (ajprincess):

sin2A=2sinAcosA cos2A=cos^2A-sin^2A=1-2sin^2A=2cos^2A-1 tan2A=2tanA/1-tan^2A. which one of these resembles the one given in ur question?

OpenStudy (anonymous):

Thank you I got - sqrt 3/2

OpenStudy (ajprincess):

welcome:)